OpenStudy (anonymous):

PLEASE HELP!! ASAP!! 1.simplify the complex fraction x/x+3/1/x+1/x+3 A. x^2/2x+3 B. x^2/2x C. x^2/x+3 D. x^2+2x+3/2x+3 2. what is the quotient in simplified form? state any restrictions on the variable

OpenStudy (anonymous):

\[\frac{ 4-x }{ x^2+5x-6 }\div \frac{ x^2-11x+28}{ x^2+7x+6 }\] A.\[\frac{ -(x+1) }{ (x-1)(x-7) }, x \neq 1, x \neq 7\] B. \[\frac{ -1 }{ (x-7) }, x \neq 7\] C. \[\frac{ -1 }{ (x-7 }, x \neq 1, x \neq -6, x \neq4\] D. \[\frac{ -(x+1) }{ (x-1)(x-7) }, x \neq1, x \neq-6, x \neq4, x \neq7\]